OpenStudy (alexatiger):

Hi guys; I need help understanding how sin x times sec x equals opposite over adjacent. I have no clue (-.-)

TheSmartOne (thesmartone):

Do you know what sin is in terms of opposite, adjacent and hypotenuse? Also: Hint: \(\large\sf sec ~x = \frac{1}{cos~x}\)

OpenStudy (alexatiger):

Sine equals opposite over hypotenuse. (Going in the right direction?)

TheSmartOne (thesmartone):

Correct! Now, what is cos = ?

OpenStudy (alexatiger):

Cosine equals adjacent over hypotenuse. :D

TheSmartOne (thesmartone):

yup! :D now, sec x = 1/cos x so to find sec in those terms, we need to find the reciprical of cos x the reciprical of a/b is b/a so what is the recirpocal of adjacent/hypotenuse?

OpenStudy (alexatiger):

Hypotenuse over adjacent

TheSmartOne (thesmartone):

correct, now sin x * sec x let's re-write them in the terms we said: \(\sf\Large \frac{opposite}{hypotenuse} \times \frac{hypotenuse}{adjacent} =\) now simplify it :)

OpenStudy (alexatiger):

opposite times adjacent? HOLY CRAP I GET IT LOL Thank you!!!

TheSmartOne (thesmartone):

Yay! Correct! Now... wait a minute. Did you mean opposite over adjacent? :P

OpenStudy (alexatiger):

No D:

OpenStudy (alexatiger):

Oh wait, since they canceled out, can't I put in 1 and then multiply and get opposite over adjacent?

TheSmartOne (thesmartone):

ok, you see: \(\sf\Large \frac{opposite}{\cancel{hypotenuse}} \times \frac{\cancel{hypotenuse}}{adjacent} =\) since we have it in the denominator and the numerator

OpenStudy (alexatiger):

XD Yes. I think I understand now :D

TheSmartOne (thesmartone):

Great! :D Also, by the way, opposite/adjacent is tan :P Just something you might want to remember. Never know when it will help. And tan = sin/cos That's all! :D

OpenStudy (alexatiger):

Wow :0 thanks for telling me!

TheSmartOne (thesmartone):